OpenStudy (anonymous):

Solve P = 2w + h for h.

OpenStudy (anonymous):

h = 1 I'm pretty sure

OpenStudy (anonymous):

off the top of my head anyway.

OpenStudy (anonymous):

okay do you know what they are asking you to do?

OpenStudy (anonymous):

Someone want to go look at the question I have there please, it's to do with graphing and applications of quadratic formulas

OpenStudy (anonymous):

@dread907 please do not just give him the answer like you already attempted to do

OpenStudy (anonymous):

i think solve for h

OpenStudy (anonymous):

yes but that means to isolate h on one side of the equation with everything else on the opposite side

OpenStudy (anonymous):

@whatisthequestion yeah I figured that wasn't right, which is why i said that and sorry about that, I need help with mine I can figure out the original area, just not what the border should be.

OpenStudy (anonymous):

ok right

OpenStudy (anonymous):

where are you guys getting numbers from all I see are variables

OpenStudy (anonymous):

and you cannot use quadratic when you have three different variables

OpenStudy (anonymous):

The variables alone can substitute automatically for 1.

OpenStudy (anonymous):

h = P + 2w i think this is right

OpenStudy (anonymous):

almost but you do not add 2w to both sides that would give you h + 4w = p + 2w

OpenStudy (anonymous):

What would take away 2w from the right hand side instead of adding another 2w?

OpenStudy (anonymous):

2w + P = H that's all they are asking I'm pretty sure

OpenStudy (anonymous):

h = P – 2w is that better

OpenStudy (anonymous):

yes @learninguy4 you have correctly solved the equation just remember to get rid of addition you subtract from both sides not add

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

@dread907 you did the same as learninguy4's second last answer please look at my response to him but where did you get that the number one could be used in place of the variables?

OpenStudy (anonymous):

@learninguy4 please remember to close the question and medal to whoever helped you the most.

OpenStudy (anonymous):

*and give a medal to

OpenStudy (anonymous):

From my algebra course, and from Geometry, a variable can be said to be as 1 because the simplicity of it, but I forgot that's if your working with one of the same variable such as 4x - x would be 3x because the variable x would act as 1

OpenStudy (anonymous):

if that works for you that is great and I will not ruin it for you but as you jut saw from the current question that explanation has some serious flaws

OpenStudy (anonymous):

Yes, that was my fault completely in an equation formed like this you cannot substitute 1 for x, only in the example above could you substitute 1 for x, and of course questions related to that. That rule is not confined to that one problem, however questions like it. So to sum up what I'm saying and whatisthequestion is getting at, then you can substitute 1 for variable so long as they are working with the same variable as the other. (That's a little confusing but that's all I can think to do to explain it sorry.)